# Improved GUI Grounding via Iterative Narrowing

### v2

@@ Abstract
Graphical User Interface (GUI) grounding plays a crucial role in enhancing the capabilities of Vision-Language Model (VLM) agents. While general VLMs, such as GPT-4V, demonstrate strong performance across various tasks, their proficiency in GUI grounding remains suboptimal. Recent studies have focused on fine-tuning these models specifically for one-shot GUI grounding, yielding significant improvements over baseline performance. We introduce a visual prompting framework that employs an iterative narrowing mechanism to improve the performance of both general and fine-tuned models in GUI grounding by up to 61%. For evaluation, we tested our method on a comprehensive benchmark comprising various UI platforms and provided the code to reproduce our results in this GitHub repository.
@@ 3 Methodology
Given an input image and a corresponding text query, our method begins by preprocessing the image to standardize its size to $9999\times 9$ 999 pixels. The model is then prompted to predict normalized coordinates $(x,y)$ that identify a relevant location on the image. Using these predicted coordinates, we create an image crop centered around the prediction for the next iteration (see Figure 1). If the original image has a landscape orientation, the new crop has its width halved and its height halved upon each iteration. For portrait-oriented images, the width is reduced to 80% of its original size and the height is reduced to one-third of the original size.
This iterative process of generating new coordinates and refining the image crop continues for $n$ iterations. On the final iteration, we take the coordinates generated as the model’s prediction of the target location after converting them to be relative to the original image.
@@ 4.2 Weaknesses & Observations
A key weakness of our method is the loss of contextual information as iterations progress, posing challenges for identifying target elements that rely on spatially distant cues. GUI-specific models like OS-Atlas-Base-7B [9] can accurately identify such targets in the baseline, while our method often fails due to limited context.
For example, in a table screenshot with text labels on the left and action buttons on the right, our method may focus on a cropped area containing only buttons ($\texttt{A}_{\texttt{button}},\texttt{B}_{\texttt{button}},\texttt{C}_{\texttt{button}}$), losing sight of the labels needed to correctly identify the target.

### v4

@@ Abstract
Graphical User Interface (GUI) grounding plays a crucial role in enhancing the capabilities of Vision-Language Model (VLM) agents. While general VLMs, such as GPT-4V, demonstrate strong performance across various tasks, their proficiency in GUI grounding remains suboptimal. Recent studies have focused on fine-tuning these models specifically for one-shot GUI grounding, yielding significant improvements over baseline performance. We introduce a visual prompting framework that employs an iterative narrowing mechanism to further improve the performance of both general and fine-tuned models in GUI grounding. For evaluation, we tested our method on a comprehensive benchmark comprising various UI platforms and provided the code to reproduce our results in this GitHub repository.
@@ 3 Methodology
Our method refines the prediction of target coordinates within an input image $\mathcal{I}$, guided by a text query $\mathcal{Q}$, through an iterative process of cropping and coordinate prediction.
Initially, the entire image $\mathcal{I}$ is treated as the cropping window, with dimensions equal to the original image size. At each iteration, the model predicts normalized coordinates $(x_{k},y_{k})\in[0,1]\times[0,1]$, representing a relevant location within the current cropping window. These coordinates are used to update the cropping window for the next iteration, progressively narrowing the region of focus.
The cropping window dimensions are adjusted based on the image orientation. For landscape-oriented images, the width and height are halved at each iteration. In the case of portrait-oriented images, the width is scaled less aggressively by a factor of $1/1.2$, while the height is halved.
After $n$ iterations, the final normalized coordinates are transformed into absolute coordinates relative to the original image dimensions. This transformation ensures that the output $(X,Y)$ corresponds directly to a location within the original image.
3.1 Pseudocode
The following pseudocode outlines the iterative process of predicting the target location in the image:
Algorithm 1 Iterative Cropping and Prediction
Input: Image $\mathcal{I}$, Query $\mathcal{Q}$, Number of iterations $n$
Output: Final coordinates $(X,Y)$ relative to the original image
  1. 1.
    Initialize cropping window $\mathcal{C}_{0}$ dimensions $w_{0}=W_{\text{orig}},h_{0}=H_{\text{orig}}$
  2. 2.
    For $k=1$ to $n$:
    1. (a)
      Predict normalized coordinates $(x_{k},y_{k})\in[0,1]\times[0,1]$ relative to $\mathcal{C}_{k-1}$, from $\mathcal{Q}$
    2. (b)
      If $\mathcal{I}$ is landscape-oriented:
      • $w_{k}\leftarrow\frac{w_{k-1}}{2},\quad h_{k}\leftarrow\frac{h_{k-1}}{2}$
    3. (c)
      Else (portrait-oriented):
      • $w_{k}\leftarrow\frac{w_{k-1}}{1.2},\quad h_{k}\leftarrow\frac{h_{k-1}}{2}$
    4. (d)
      Define cropping window $\mathcal{C}_{k}$:
      • Center: $(x_{k},y_{k})$
      • Dimensions: $(w_{k},h_{k})$
  3. 3.
    Convert normalized $(x_{n},y_{n})$ to coordinates relative to the original image space $(X,Y)$
  4. 4.
    Return Final coordinates $(X,Y)$
@@ 4.2 Weaknesses & Observations
A key weakness of our method is the loss of contextual information as iterations progress, posing challenges for identifying target elements that rely on spatially distant cues. GUI-specific models like OS-Atlas-Base-7B [9] can accurately identify such targets in the baseline, while our method often fails due to limited context (see Figure 2).